OpenStudy (anonymous):

simplify: (16y^8/4y^2)^-4

OpenStudy (anonymous):

\[\frac{ 1 }{ 256y^{40} }\]

OpenStudy (andriod09):

so what do you already know how to do on this question?

OpenStudy (anonymous):

nothing. im completely confused.

OpenStudy (andriod09):

wait, how do you have an exponent on an exponent?

OpenStudy (anonymous):

@Corey_ONeal I tried that answer it wasnt right.

OpenStudy (anonymous):

what do you mean?

OpenStudy (andriod09):

\[(16y^8\div4y^2)^{-4}\] nvm. i found it.

OpenStudy (anonymous):

ok

OpenStudy (andriod09):

so what do you know about it?

OpenStudy (anonymous):

im not really sure. im helping a friend out and now im frustrated because i cant figure out the answer to the question.

OpenStudy (anonymous):

do you know how to do it?

OpenStudy (andriod09):

use order of operations to help you. P arenthesis E xponents M ultiplication D ivision A ddition S ubtraction

OpenStudy (anonymous):

im still really confused. i did the parentheses which is 16/4=4. then exponents is 8/2=4.

OpenStudy (andriod09):

you don't divide exponets, you multiply them by the product of the answer. and it would be: \[16y^8\div4y^2=4^2y^2\] which would give you \[4^2y^2=16y^2\]

OpenStudy (andriod09):

\[16y^4\] my bad

OpenStudy (anonymous):

what do i do after that?